# Fisher-Geometric Diffusion in Stochastic Gradient Descent: Optimal Rates, Oracle Complexity, and Information-Theoretic Limits

### 02 Mar 2026 v1

@@ 9 Discussion
A recurring theme in the empirical “folklore” around SGD is that batch size, stepsize, and curvature interact in ways that are not captured by treating the stochastic gradient as an arbitrary noisy oracle. The main message of this paper is that the noise is not arbitrary: under standard mini-batch sampling it inherits a model-determined covariance, and that covariance induces the natural statistical geometry for both analysis and design. In the diffusion/OU regime, SGD is therefore best viewed as a controlled stochastic dynamical system whose steady-state and finite-time behavior are governed by a Lyapunov balance rooted in the intrinsic statistical geometry.
@@ 9.1 Summary of Contributions
We derive the leading-order mini-batch covariance directly from sampling design and exchangeability: the batch-mean covariance is a finite-population correction factor times an intrinsic gradient-noise covariance. For likelihood losses under correct specification this intrinsic object is the (projected) Fisher information; for general losses it is the projected gradient covariance (and the associated sandwich/Godambe geometry). This identifies the diffusion-matrix in SDE approximations of SGD at the level of sampling theory rather than postulation.
Once the intrinsic geometry is identified, the right performance metric follows. We prove a Fisher/Godambe risk upper bound with the canonical $1/T$ scaling (per iteration) and match it with an information-theoretic lower bound (per oracle call). The result isolates (i) an intrinsic dimension $d_{\operatorname{eff}}$ rather than ambient $d$, and (ii) statistical identifiability through the relevant nondegeneracy of the intrinsic statistical geometry, as the fundamental limiters of stochastic optimization speed.
Building on these risk bounds, we translate the theory into a high-probability oracle-complexity statement for $\varepsilon$-stationarity in the Fisher dual norm. The resulting scaling is governed by a Fisher/Godambe condition number, not the Euclidean/Hessian condition number that dominates deterministic analyses. This clarifies why problems can be Euclidean-stiff yet statistically well-conditioned, and why SGD can be efficient in such settings.
Scope of results: proven, sketched, and deferred.
To help the reader calibrate what is fully established here versus what is indicated:
  • Proven in this paper. Alignment (Theorem 4.3), diffusion approximation (Theorem 5.4), OU/Lyapunov balance (Corollary 5.8), matching upper–lower bounds in the Fisher metric (Theorems 6.46.8), and oracle complexity (Theorem 7.3). All hold under the local regularity conditions of Section 3.
  • Sketched / partially treated. The batch-size scaling corollary (Corollary 7.8) is a direct consequence of the complexity bound but does not optimize over non-constant batch schedules. The variance-reduction bridge (Remark 9.1) is stated for uniform Loewner reductions; direction-specific VR requires further analysis.
  • Deferred to a companion paper. Extensions to model misspecification (sandwich geometry at the KL projection), weak dependence (long-run variance corrections), explicit batching policies via covariance control, and OR-specific applications (simulation optimization, revenue management, inventory). Appendix B provides preliminary experiments along the misspecification and approximate-exchangeability directions.
@@ 9.2 Implications for Algorithm Design
In the diffusion regime the key knob is the effective temperature $\tau=\eta/b$. Changing $b$ does more than “reduce variance”; it changes the diffusion amplitude in a directional way determined by Fisher/Godambe geometry, and therefore changes the Lyapunov-predicted error floor. This makes precise the heuristic that small batches promote exploration: they inject correctly shaped anisotropic noise that preferentially excites statistically flat directions.
Because the intrinsic noise is matrix-valued, variance-reduction interventions (control variates, antithetics, CRN, stratification, etc.) should be evaluated by their effect on risk in a chosen metric (e.g., $\mathrm{Tr}(Q\Sigma)$ or the induced $\mathrm{Tr}(QM^{\star})$ terms) rather than Euclidean variance alone. In this way, classical OR variance-reduction practice connects to SGD through the correct statistical geometry.
@@
A classical VR technique replaces $G^{\star}$ with a reduced covariance $\widetilde{G}\preceq G^{\star}$ in the Loewner order. Through the Lyapunov balance (Corollary 5.8), this yields a reduced stationary covariance $\widetilde{\Sigma}_{\infty}\preceq\Sigma_{\infty}$ and hence lower Fisher-metric risk: $\mathrm{Tr}(F^{\star}\widetilde{\Sigma}_{\infty})\leq\mathrm{Tr}(F^{\star}\Sigma_{\infty})$. Thus any VR method that uniformly reduces the gradient-noise covariance also reduces the Fisher-weighted risk, providing a direct bridge between classical simulation efficiency and our geometric framework.
The Fisher/Godambe object shapes fluctuations (diffusion) and thus enters rates and stationary covariance, but SGD does not explicitly precondition the drift. Natural-gradient methods act on the drift by reweighting $\nabla L$, whereas our results explain why plain SGD exhibits implicitly geometry-aware behavior through its noise structure. A unified view treats drift contraction and diffusion injection as distinct levers.
These observations lead to a direct route to adaptive batching. The control view suggests increasing batch size (cooling temperature) when the marginal value of variance reduction is large in the performance metric of interest, and decreasing it when exploration is valuable. This provides a principled alternative to hardware-only or trial-and-error batch schedules, and it clarifies the complementary roles of stepsize (contraction) and batch size (temperature).
@@ 9.3 Practical Notes and Extensions
Frozen-metric regime (“basin of Fisher geometry”).
Our analysis freezes $F(\theta)$ (or $G(\theta)$) locally at $F^{\star}$, a standard linearization in diffusion/OU approximations. A practical way to assess the validity of this regime is to monitor the relative drift of the metric along the trajectory:
$\delta_{F}(t):=\frac{\|F(\theta_{t})-F^{\star}\|_{\operatorname{op}}}{\|F^{\star}\|_{\operatorname{op}}},\qquad\delta_{H}(t):=\frac{\|H(\theta_{t})-H^{\star}\|_{\operatorname{op}}}{\|H^{\star}\|_{\operatorname{op}}}.$
In our experiments (Section 8), the Lyapunov predictions remain accurate even from random starts, suggesting that the algorithm enters a basin of Fisher geometry early; quantitatively, this corresponds to epochs where $\delta_{F}(t)$ and $\delta_{H}(t)$ stay modest (single-digit to tens of percent).
Estimating $G^{\star}$ online.
In simulation-optimization settings where $G^{\star}=\mathrm{Var}(\psi(\theta^{\star};Z))$ is not available analytically, one can estimate it on the fly using standard streaming covariance updates. If per-sample gradients are available within a mini-batch, a running (or exponentially weighted) covariance of $\nabla\ell(\theta_{t};Z_{i})$ is consistent under stationarity; when only batch gradients are available, one may estimate $\mathrm{Var}(g_{t})$ across iterations and recover an implied per-sample covariance up to the $1/b$ scaling. Importantly, many of our design implications depend only on low-order summaries (e.g., $\mathrm{Tr}(G^{\star})$ and leading eigendirections), which can be tracked cheaply via trace estimators and low-rank streaming PCA.
Adaptive batch sizing as temperature control.
Since $\tau=\eta/b$ acts as an effective temperature in the OU approximation, a natural adaptive strategy is to regulate $\tau_{t}$ based on a local curvature proxy. For example, one may choose $b_{t}$ (or $\eta_{t}$) to keep $\tau_{t}$ approximately proportional to $1/\lambda_{\max}(H(\theta_{t}))$ (or a trace/diagonal surrogate), thereby preventing over-heating in stiff directions while avoiding unnecessary sampling when curvature is mild. We view such schedules as practical corollaries of the diffusion model: the analysis identifies the control variable ($\tau$) and the geometric quantity it should track (local curvature), while leaving the controller design to the application.
High-dimensional and over-parameterized regimes.
When $d_{\operatorname{eff}}\ll d$, the analysis suggests that the relevant geometry is concentrated in an identifiable subspace captured by the projection $\Pi_{\theta}$ (Definition 3.5). In this regime, plain SGD can exhibit behavior reminiscent of natural-gradient dynamics within the Fisher-active subspace, while remaining insensitive to unidentifiable directions that carry little Fisher information. This perspective aligns with the empirical observation that diffusion/Lyapunov predictions often remain accurate in large models: what matters is not the ambient dimension $d$, but the effective dimension $d_{\operatorname{eff}}$ induced by the Fisher/Godambe structure.
Summary. Taken together, these notes reinforce the central message: the theory identifies a small set of geometric primitives—local curvature $H$, noise covariance $G$, and effective temperature $\tau=\eta/b$—that govern both transient rates and stationary risk. This makes the framework naturally compatible with practice: the primitives can be monitored online (even approximately), and algorithmic knobs (learning rate and batch size) can be interpreted as temperature-control mechanisms that respect the identifiable Fisher geometry.
@@ 9.4 Limitations and Future Work
Our sharpest statements use local linearization near a nondegenerate critical point. Extending the theory to genuinely nonconvex regimes requires controlling state-dependent curvature and noise and relating the intrinsic statistical geometry to metastability and basin transitions.
When $d\gg n$ or models are overparameterized, Fisher/Godambe objects may be low-rank. A fully satisfactory theory in this regime should work intrinsically on identifiable manifolds (or with pseudoinverses/regularized metrics), and it should clarify how to define $d_{\operatorname{eff}}$ and stationarity when the geometry is degenerate.
Remark 9.2 (Practical workaround for low-rank Fisher/Godambe).
In the overparameterized regime where $\mathrm{rank}(F^{\star})=r\ll d$, the projection $\Pi_{\theta^{\star}}$ onto the identifiable tangent space (Definition 3.5) restricts all objects to an $r$-dimensional subspace. The theory then applies with $d$ replaced by $r$ throughout, and $d_{\operatorname{eff}}$ is computed from the $r$ nonzero eigenvalues of $F^{\star}$. In practice, this amounts to working in the span of the top eigenvectors of an empirical Fisher/Godambe estimate, which is a standard preprocessing step in natural-gradient and KFAC implementations.
Methods such as Gauss–Newton, diagonal preconditioners, and KFAC reshape the drift and can also change the effective noise geometry through the induced mini-batch estimators. A joint control theory that treats drift contraction and diffusion covariance as co-designed objects would connect information geometry to OR-style variance-allocation at scale.
In distributed settings, batch size trades off with communication and client heterogeneity. Incorporating these constraints into the covariance recursion and deriving optimal schedules under explicit communication budgets is a natural extension of the OR control formulation developed here.
Key Results Summary
Table 5: Key results summary.
Result Equation
Godambe/Fisher alignment $\mathrm{Cov}\!\bigl(g_{B}(\theta^{\star})\bigr)=\frac{1}{b}\,G^{\star}(\theta^{\star})$
OU / diffusion approximation $d\Delta_{s}=-H^{\star}\Delta_{s}\,ds+\sqrt{\tau}\,(G^{\star})^{1/2}dW_{s},\;\;\tau=\eta/b$
Lyapunov equilibrium $H^{\star}\Sigma_{\infty}+\Sigma_{\infty}{H^{\star}}^{\!\top}=\tau\,G^{\star}(\theta^{\star})$
Upper bound (frozen metric) $\mathbb{E}\!\big[\|\theta_{T}-\theta^{\star}\|_{F^{\star}}^{2}\big]\lesssim\frac{1}{Tb}$
Lower bound (minimax) $\inf\sup\,\mathbb{E}\!\big[\|\hat{\theta}_{T}-\theta^{\star}\|_{F^{\star}}^{2}\big]\gtrsim\frac{d}{Tb}$
Oracle complexity ($N=Tb$) $N=\Theta\!\left(\frac{\kappa_{F}\,d_{\mathrm{eff}}(F^{\star})}{\varepsilon^{2}}\log\!\frac{1}{\delta}\right)$
@@ A.1 Finite-population covariance identity
$\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}\left(1-\frac{b}{n}\right)\frac{n}{n-1}\,S_{n}=\frac{1}{b}\left(1-\frac{b-1}{n-1}\right)S_{n}.$ (48)
@@
Write $\bar{x}_{B}=\sum_{i=1}^{n}w_{i}x_{i}$ where $w_{i}=\frac{1}{b}\mathbf{1}\{i\in B\}$. Under simple random sampling without replacement of size $b$ from $\{1,\ldots,n\}$, let $I_{i}=\mathbf{1}\{i\in B\}$. The indicator moments are:
$\mathbb{E}[I_{i}]$ $=\frac{b}{n},$
$\mathrm{Var}(I_{i})$ $=\frac{b}{n}\left(1-\frac{b}{n}\right),$
$\mathrm{Cov}(I_{i},I_{j})$ $=-\frac{b(n-b)}{n^{2}(n-1)}\quad\text{for }i\neq j.$
Since $w_{i}=\frac{1}{b}I_{i}$, one computes $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\sum_{i,j}\mathrm{Cov}(w_{i},w_{j})\,x_{i}x_{j}^{\top}$. Expanding around $\bar{x}_{n}$ and using $\sum_{i}(x_{i}-\bar{x}_{n})=0$, standard algebra yields $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}(1-\frac{b}{n})\frac{n}{n-1}\,S_{n}$. The equivalence of the two forms follows from $(1-\frac{b}{n})\frac{n}{n-1}=\frac{n-b}{n-1}=1-\frac{b-1}{n-1}$. ∎
Remark A.2 (Reconciliation with main text).
The main text Lemma 4.1 states the result as $\mathrm{Cov}(g_{B}(\theta)\mid z_{1:n})=\frac{1}{b}(1-\frac{b}{n})S_{n}(\theta)$, which implicitly absorbs the factor $\frac{n}{n-1}$ into the definition of $S_{n}$. Both forms are equivalent; the main text uses the convention that $S_{n}$ denotes the population-style covariance with $1/n$ normalization, while this appendix uses the sample covariance with 1/(n1)1/(n-1) normalization and shows the correction factor explicitly.
@@ A.2 Score identities and information equality
(i) Differentiate the normalization constraint: $0=\frac{\partial}{\partial\theta_{j}}\int p_{\theta}(z)dz=\mathbb{E}_{\theta}[s_{\theta}(Z)_{j}]$.
(ii) Differentiate (i) again and use $\frac{\partial}{\partial\theta_{k}}\mathbb{E}_{\theta}[f(Z)]=\mathbb{E}_{\theta}[f(Z)\cdot s_{\theta}(Z)_{k}]$ to obtain $0=\mathbb{E}_{\theta}[\partial^{2}_{jk}\log p_{\theta}(Z)]+\mathbb{E}_{\theta}[s_{\theta}(Z)_{j}\cdot s_{\theta}(Z)_{k}]$. Rearranging gives the information equality. ∎
@@ A.3 Lyapunov equation solution
Remark A.5 (Sign convention).
In the OU process $d\Delta_{s}=-H^{\star}\Delta_{s}ds+\ldots$, the drift matrix is $H^{\star}$. For ergodicity, we need $H^{\star}$ to have eigenvalues with negative real part, which holds since $H^{\star}=\nabla^{2}L(\theta^{\star})\succ 0$ by assumption. The Lyapunov equation $H^{\star}\Sigma+\Sigma H^{\star\top}=Q$ applies Lemma A.4 with $A=H^{\star}$.
@@ Appendix B Robustness Beyond Pure Exchangeability
The experiments in this appendix demonstrate that the Fisher/Godambe alignment theory degrades gracefully under realistic departures from i.i.d. likelihood assumptions. Experiment A considers model misspecification, where curvature and noise geometry diverge; Experiment B models mild violations of exchangeability via variance inflation.
@@ B.1 Experiment A: Misspecification (Probit DGP, Logistic Fit)
Setup.
We generate data from a probit DGP and fit a logistic model by SGD, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$. Under misspecification, the score/gradient covariance and curvature no longer coincide:
$J^{\star}:=\mathrm{Var}\!\Bigl(\nabla_{\beta}\ell(\beta;X,Y)\big|_{\beta=\beta^{\star}}\Bigr),\qquad H^{\star}:=\nabla^{2}\mathbb{E}\bigl[-\ell(\beta;X,Y)\bigr]\big|_{\beta=\beta^{\star}},$
$J^{\star}\neq H^{\star}\ \text{in general}.$
Benchmark.
The local diffusion/OU surrogate predicts a sandwich-form Lyapunov balance:
$H^{\star}\Sigma^{\star}+\Sigma^{\star}H^{\star\top}=\frac{2}{b}\,J^{\star},$
which is the dynamical analogue of the classical M-estimation sandwich covariance. We compute $H^{\star}$ and $J^{\star}$ at $\beta^{\star}$ and solve the corresponding Lyapunov equation numerically.
What the figure shows.
Figure 9 shows that the misspecified theory using $J^{\star}$ tracks the empirical stationary covariance of the SGD iterates, while the naive well-specified surrogate that replaces $J^{\star}$ by $H^{\star}$ exhibits systematic bias. This validates the paper’s core message in the simplest non-ideal setting: under misspecification, curvature alone is not enough—the noise geometry is the first-class object.
Figure 9: Experiment A: misspecification. Sandwich/Lyapunov prediction (using $J^{\star}$) vs. naive $J^{\star}\!=\!H^{\star}$ surrogate and empirical stationary covariance.

### 03 Apr 2026 v2

@@ 9.1 Summary of Contributions
The paper’s central contribution is the identification of the mini-batch gradient covariance from sampling design and exchangeability as the projected gradient-noise matrix $G^{\star}(\theta)$—which reduces to Fisher information at $\theta^{\star}$ under correct specification and, for general losses, induces Godambe/sandwich geometry together with $H^{\star}$. This identification endogenizes the diffusion coefficient, determines the Lyapunov stationary covariance, and selects the natural metric for rates and complexity, yielding matching $1/N$ bounds governed by intrinsic dimension $d_{\mathrm{eff}}$ and condition number $\kappa_{F}$.
  • Proven: Alignment (Theorem 4.3), diffusion (Theorem 5.4), OU/Lyapunov (Corollary 5.7), discrete bridge (Proposition 5.9), matching bounds (Theorem 6.5 and Proposition 6.10), and oracle complexity (Theorem 7.3).
  • Sketched: Batch-size scaling (Corollary 7.8) and variance-reduction bridge (Remark 9.1).
  • Deferred: Misspecification, weak dependence, explicit batching policies, and OR-specific applications (Appendix B provides preliminary experiments).
@@ 9.2 Implications for Algorithm Design
In the diffusion regime the key knob is the effective temperature $\tau=\eta/b$. Changing $b$ does more than “reduce variance”; it changes the diffusion amplitude in a directional way determined by the identified geometry, and therefore changes the Lyapunov-predicted error floor. Because the intrinsic noise is matrix-valued, variance-reduction interventions should be evaluated by their effect on risk in a chosen metric (e.g., $\mathrm{Tr}(Q\Sigma)$) rather than Euclidean variance alone, connecting classical OR variance-reduction practice to SGD through the correct geometry.
@@
A VR technique that reduces $G^{\star}$ to $\widetilde{G}\preceq G^{\star}$ in the Loewner order yields $\widetilde{\Sigma}_{\infty}\preceq\Sigma_{\infty}$ via the Lyapunov balance, and hence lower Fisher-metric risk.
SGD does not precondition the drift; natural-gradient methods do so by reweighting $\nabla L$. Our results explain why plain SGD nevertheless exhibits geometry-aware behavior through its noise structure: drift contraction and diffusion injection are distinct levers. The control view also suggests adaptive batching—increasing $b$ (cooling) when marginal variance reduction is valuable in the performance metric of interest, and decreasing it when exploration matters.
@@ 9.3 Practical Notes and Extensions
Our analysis freezes $F(\theta)$ locally at $F^{\star}$. Validity can be assessed by monitoring $\delta_{F}(t):=\|F(\theta_{t})-F^{\star}\|_{\operatorname{op}}/\|F^{\star}\|_{\operatorname{op}}$; in our experiments the Lyapunov predictions remain accurate from random starts, indicating early entry into the basin of Fisher geometry.
9.4 Estimating $G^{\star}$ in Practice
The preceding theory identifies $G^{\star}(\theta)$ as the central primitive. In applications where $G^{\star}$ is not available analytically, several estimation strategies provide operationally useful approximations. The simplest is the empirical covariance of per-sample gradients: given a pilot batch $\{z_{1},\ldots,z_{m}\}$ at $\theta_{t}$, the sample covariance $\hat{G}(\theta_{t}):=\frac{1}{m-1}\sum_{i=1}^{m}(\psi_{i}-\bar{\psi})(\psi_{i}-\bar{\psi})^{\top}$ (where $\psi_{i}:=\nabla_{\theta}\ell(\theta_{t};z_{i})$) is the natural plug-in estimator and converges at the standard $O(1/\sqrt{m})$ rate in operator norm. For large $d$, a diagonal approximation $\hat{G}_{\mathrm{diag}}$ retains the per-coordinate noise scales and suffices for estimating $d_{\mathrm{eff}}$ and detecting gross anisotropy; when the leading eigendirections of $G^{\star}$ are the primary target, a rank-$k$ approximation via streaming PCA on per-sample gradients requires only $O(dk)$ storage. If $G^{\star}(\theta)$ varies slowly along the optimization path, an exponential moving average $\hat{G}_{t}=(1-\alpha)\hat{G}_{t-1}+\alpha\,(\psi_{t}-\bar{\psi}_{t})(\psi_{t}-\bar{\psi}_{t})^{\top}$ tracks the local geometry with minimal overhead.
[…]
@@ 9.5 Limitations and Future Work
Our sharpest statements use local linearization near a nondegenerate critical point; extending to nonconvex regimes requires controlling state-dependent curvature and relating intrinsic geometry to metastability. When $d\gg n$, the identified covariance may be low-rank, and a satisfactory theory should work on identifiable manifolds with appropriate pseudoinverses. Finally, preconditioners (Gauss–Newton, KFAC) reshape both drift and noise geometry; a joint control theory co-designing contraction and diffusion—potentially under communication constraints—would connect information geometry to OR-style variance allocation at scale.
@@ A.1 Finite-population covariance identity
$\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=\frac{1}{b}\left(1-\frac{b}{n}\right)S_{n}.$ (50)
@@
Write $\bar{x}_{B}=\frac{1}{b}\sum_{i=1}^{n}I_{i}x_{i}$ where $I_{i}=\mathbf{1}\{i\in B\}$. Under SRSWOR, $\mathbb{E}[I_{i}]=b/n$, $\mathrm{Var}(I_{i})=\frac{b}{n}(1-\frac{b}{n})$, and $\mathrm{Cov}(I_{i},I_{j})=-\frac{b(n-b)}{n^{2}(n-1)}$ for $i\neq j$. Expanding $\mathrm{Cov}(\bar{x}_{B}\mid x_{1:n})=b^{-2}\sum_{i,j}\mathrm{Cov}(I_{i},I_{j})x_{i}x_{j}^{\top}$ and collecting terms:
$b^{-2}\!\left[\frac{b(n-b)}{n^{2}}\sum_{i}x_{i}x_{i}^{\top}-\frac{b(n-b)}{n^{2}(n-1)}\sum_{i\neq j}x_{i}x_{j}^{\top}\right]$
$\quad=\frac{n-b}{bn^{2}}\cdot\frac{n}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}=\frac{1}{b}\!\left(1-\frac{b}{n}\right)S_{n},$
where the last equality uses $(n-1)S_{n}=\sum(x_{i}-\bar{x}_{n})(x_{i}-\bar{x}_{n})^{\top}$. ∎
@@ A.2 Score identities and information equality
Standard; see, e.g., Lehmann and Casella (1998, Ch. 7). ∎
@@ Appendix B Robustness Beyond Pure Exchangeability
The experiments in this appendix are exploratory robustness diagnostics, not part of the formal theoremic validation in Section 8. They demonstrate that the covariance-identification theory degrades gracefully under realistic departures from i.i.d. likelihood assumptions. Experiment A considers model misspecification, where curvature and noise geometry diverge; Experiment B models mild violations of exchangeability via variance inflation.
@@ B.1 Experiment A: Misspecification (Probit DGP, Logistic Fit)
We generate n=50,000n=50{,}000 observations from a probit DGP and fit a logistic model, so that the optimization target is the KL projection $\beta^{\star}$ rather than the true $\beta_{0}$. Under misspecification the score covariance and curvature no longer coincide: $J^{\star}:=\mathrm{Var}\!\bigl(\nabla_{\beta}\ell(\beta;X,Y)\big|_{\beta=\beta^{\star}}\bigr)\neq H^{\star}:=\nabla^{2}\mathbb{E}[-\ell(\beta;X,Y)]\big|_{\beta=\beta^{\star}}$. The covariance-identification theorem (Theorem 4.3) nevertheless predicts $\mathrm{Cov}(g_{B}(\beta^{\star}))\approx(1/b)\,J^{\star}$, where $J^{\star}$ plays the role of $G^{\star}(\theta^{\star})$ in the general-loss notation. We estimate $J^{\star}$ from the full-sample scores and compare its diagonal entries to those of the empirical mini-batch covariance (batch size $b=256$, 800 replicates). Figure 9 shows close agreement (relative Frobenius difference $\approx 0.05$), confirming that the identification remains accurate under misspecification: mini-batch fluctuations are governed by the score covariance $J^{\star}$ rather than by curvature alone.
Figure 9: Experiment A: misspecification (probit DGP, logistic fit). Diagonal entries of the empirical mini-batch covariance vs. $(1/b)\,\hat{J}^{\star}$ (projected score covariance at $\beta^{\star}$); relative Frobenius difference $\approx 0.05$.